At least how many Calories does a mountain climber need in order to climb from sea level to the top of a 5.07 km tall peak assuming the muscles of the climber can convert chemical energy to mechanical energy with an efficiency of 16.0 percent. The total mass of the climber and the equipment is 82.0 kg. (Enter your answer as a number without units.)
Answer #1

given that

mass m = 82 kg

height h = 5.07*1000 = 5070 m

from the relation

gravitational potential energy U = mgh

U = 82*9.8*5070

U = 4074252 J

converting joules to calories

U = 4074252/4.2*1000 = 970.06 cal

but efficiency is 16%

U = 970.06/0.16 = 6062.875 cal